find a number whose product with 9 is the same as its sum with 65

Answer by Maths68(1474)

Let number = x
Product with 9 = 9x
Sum with 65 = x+65
According to the condition
9x=x+65
9x-x=65
8x=65
x=65/8
x=8 1/8 or 8.125